Two cars A and B start from a given point and travel on a straight road at average speeds of 70 km/h and 90 km/h respectively. If B starts 1 h after A find out how long before A traveled before B overtook it and the distance each traveled before they met.
D=R.T
d=70t d=90t
d=70t+90t
d=160t
d/160=t

Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
ok. Looks like you gave it a shot!!! I am assuming that the cars are going in the same direction. If they were going toward each other, we would have to know how far apart they were to begin with.
When car B starts, car A will have already travelled 70*1=70km.
Car B will have overtaken car A when they have both travelled the same distance
Let t =the amount of time needed for them to travel the same distance AFTER car B starts.
Distance travelled by car A=70+70t
Distance travelled by car B=90t
When these distances are the same, the two cars are together, soooo:
70+70t=90t
20t=70
t=3.5 hr
Distance car A travelled =70+70*3.5=315km
Distance car B travelled=90*3.5=315km

Hope this helps----ptaylor